# Information Extraction from Clinical Notes: Are We Ready to Switch to Large Language Models?

### v3

@@ Methods:
We investigated two fundamental clinical IE tasks: Named Entity Recognition (NER) and Relation Extraction (RE). Using 1,588 clinical notes from four data sources including UT Physicians (UTP), MTSamples, MIMIC-III, and i2b2, we developed a comprehensive annotated corpus covering 4 main clinical entities (problems, tests, medications, and other treatments) and 16 modifiers (e.g., negation, certainty). We instruction-tuned LLaMA-2 and LLaMA-3 for both tasks, comparing their performance, generalizability, computational resources, and throughput to BiomedBERT.
@@ Results:
LLaMA models consistently outperformed BiomedBERT across datasets and tasks. With sufficient training data (e.g., UTP), LLaMA showed marginal improvements (1% on NER, 1.5-3.7% on RE) over BiomedBERT; for limited training data (e.g., MTSamples, MIMIC-III), the improvements were greater. On the unseen i2b2 dataset, LLaMA-3-70B outperformed BiomedBERT by over 7% (F1) on NER and 4% (F1) on RE. However, LLaMA models required significantly more computing resources (i.e., memory and GPUs) and ran up to 28 times slower than BiomedBERT. To facilitate LLM use in clinical IE, we implemented ”Kiwi,” a clinical IE package featuring both BiomedBERT and LLaMA models and made it available at https://kiwi.clinicalnlp.org/.
@@ Conclusion:
This study is among the first to develop and evaluate a comprehensive clinical IE system using open-source LLMs. Results indicate that LLaMA models outperform BiomedBERT for clinical NER and RE but with higher computational costs and lower throughputs. These findings highlight that choosing between LLMs and traditional deep learning methods for clinical IE applications should remain task-specific, taking into account both performance metrics and practical considerations such as available computing resources and the intended use case scenarios.
@@ 1 Introduction
In response to these challenges, we developed a comprehensive clinical IE corpus, covering major clinical entities, using 1,588 clinical notes from various institutions. This corpus includes 4 main entities and their 16 corresponding modifiers. Using our corpus, we instruction-tuned and evaluated LLaMA-2 and LLaMA-3 models for both NER and RE. We systematically compared these models with BiomedBERT in terms of performance, generalizability, throughput, and energy consumption to provide insights for users on selecting the right tool for clinical IE. To facilitate real-world application, we introduced “Kiwi”, a ready-to-use clinical IE pipeline that includes our instruction-tuned LLaMA-based and BiomedBERT-based models. An online demo is available at: https://kiwi.clinicalnlp.org/.
@@ 2.1 Task
Our study includes two clinical IE tasks: Named Entity Recognition (NER) and Relation Extraction (RE). NER identifies entity types and boundaries in text 33, while RE identifies relationships between entity pairs 34. We instruction-tuned LLaMA-2 and LLaMA-3 models and compared them to fine-tuned BiomedBERT. We focused on extracting four main entities—Medical Problems, Tests, Drugs and other Treatments—along with 16 modifiers (e.g., Severity, Temporal, Dosage). In total, there are 21 possible relations between main entities and modifiers, as detailed in Table 1. Annotation methods and guidelines are available in the Supplementary File and GitHub repository.
@@ 2.3.2 Bidirectional Encoder Representations from Transformers (BERT)
We evaluated BiomedBERT 41 for NER and RE tasks. BiomedBERT was selected over other clinical domain BERT models due to its better performance observed in our preliminary experiments. NER settings included a learning rate of 5e-5, 20 epochs, weight decay of 0.01, and batch size of 4. For RE, we used a learning rate of 1e-6, 20 epochs, and batch size of 64.
@@ 2.5.1 Performance and significance test
We evaluate RE and NER performance using Precision (P), Recall (R), and F1 scores under exact and relaxed match criteria. Exact matches require entity types, boundaries, and relations to align perfectly, while relaxed matches require only entity type and relation matches with overlapping boundaries. Significance tests assess improvements over BiomedBERT, calculating p-values via a one-sided Wilcoxon rank-sum test with bootstrapping (sample size equal to test size, 1000 repetitions, 95% confidence interval).
@@ 2.5.4 Throughput
We assess each model’s throughput during fine-tuning and inference. For fine-tuning, we measure the total time to train and calculate seconds per note by dividing the total time by the number of notes in the dataset. For inference, we determine the time per note using the test dataset with a batch size of 100 for both LLaMA and BiomedBERT models.
@@ 3.1 Fine-tuning performance
Table 2 highlights the exact match performance of LLaMA-2, LLaMA-3, and BiomedBERT models, with full metrics in Supplementary Table S2. LLaMA models consistently outperformed BiomedBERT, particularly with limited fine-tuning data. With sufficient data (e.g., UTP dataset), LLaMA models show minor improvements of  1% on NER and 1.5-3.7% on RE. However, with less data (e.g., MTSamples, MIMIC-III), LLaMA models show greater performance improvement compared to BiomedBERT, with 3.7-4.5% higher F1 scores on NER and up to 12.5% higher on RE.
@@
BiomedBERT 0.898 0.670 0.808 0.703
Table 2: Exact F1 scores of LLaMA-2, LLaMA-3, and BiomedBERT on NER and RE across three datasets and cross-institution generalizability test results on unseen dataset (i2b2). All results of LLaMA models are statistically significant compared to BiomedBERT based on the one-sided Wilcoxon rank-sum test.
To explore the performance of LLaMA and BiomedBERT in low-resource settings, we conducted dataset-specific fine-tuning by training two models—LLaMA-3-70B and BiomedBERT—separately on the MTSamples and MIMIC-III datasets. Each model was fine-tuned exclusively on either MTSamples or MIMIC-III, then evaluated across all four test sets. Based on the results in Table 2, where LLaMA-3-70B demonstrated superior performance across datasets, we prioritized using this model for these experiments. Our findings show that LLaMA-3-70B consistently outperformed BiomedBERT across datasets in these low-resource scenarios, as shown in Supplementary Tables S3 and S4.
@@ 3.3 Computational resources and throughput
We compared the computational resources and throughput of BiomedBERT, LLaMA-2, and LLaMA-3 using NVIDIA A100 80GB GPUs (Tables 3 and 4). Fine-tuning was conducted with 1 GPU for BiomedBERT and the LLaMA. For inference, BiomedBERT and the 7B, 8B, and 13B LLaMA models used 1 GPU, whereas the 70B models required 2 GPUs. Results show notable trade-offs between performance and computational efficiency across model sizes.
@@ 3.3.1 Memory usage
Memory requirements increase substantially with model size. The 70B versions of LLaMA-2 and LLaMA-3 required 58-66 GB for training and 146 GB for inference, while BiomedBERT used only 21.3 GB and 65.7 GB, respectively. The 7B and 8B LLaMA models showed comparable memory efficiency to BiomedBERT, with training memory usage between 12.7-32.4 GB. Notably, during training, LLaMA models utilize Low-Rank Adaptation (LoRA), which reduces memory requirements compared to inference, where LoRA is not applied. This difference makes the memory needed for fine-tuning significantly lower than that required for inference. Additionally, these memory usages are heavily influenced by batch sizes; the current setup is optimized for best model performance.
@@ 3.3.2 Energy consumption and carbon emission
Energy consumption varied widely across models, with larger models requiring substantially more resources. LLaMA-70B models were the most energy-intensive during training, using 101-125 kWh and producing 39-49 kg of CO2). Smaller LLaMA models and BiomedBERT were more efficient, using 13-28 kWh and 5-11 kg of CO2). For inference, BiomedBERT was highly efficient at 0.03 kWh, while LLaMA models ranged from 0.14 kWh (LLaMA-3-8B) to 0.86 kWh (LLaMA-2-70B).
@@ 3.3.3 Training and Inference Throughput
Training and inference throughputs varied substantially across models and datasets as shown in table 4. LLaMA-3-8B had the fastest training time at 38.6 GPU hours, while BiomedBERT required 96.8 hours. BiomedBERT’s longer training time can be attributed to its need for 20 training epochs, compared to only 2 epochs for LLaMA models. This difference is reflected in GPU hours per epoch: BiomedBERT stands out with a particularly low requirement of 4.84 GPU hours per epoch, whereas LLaMA2-7B and LLaMA3-8B require 20.85 and 19.3, respectively. However, LLaMA-70B models took much longer, requiring 273.5-320.2 hours. Dataset characteristics such as note length impacted the training time; for example, MIMIC-III (2,106 tokens) took significantly more time per note than MTSamples (603 tokens). BiomedBERT outperformed all LLaMA models in inference, processing notes 5-28 times faster. Among LLaMA models, LLaMA-3-8B was quickest, while the 70B models were slowest. These findings highlight a trade-off: larger LLaMA models offer higher accuracy but are slower, while BiomedBERT is more efficient for time-sensitive applications.
@@ 3.4 Error analysis
Automatic and manual error analyses on LLaMA and BiomedBERT outputs across four datasets revealed that LLaMA-70B models have fewer False Negatives (FN) but slightly higher False Positives (FP) than smaller models, indicating a precision-recall trade-off with model size. BiomedBERT, while showing higher FN and FP rates, has fewer boundary errors, suggesting better accuracy in detecting text boundaries. Type errors were consistently low for all models, confirming effective classification once detected. Manual analysis of 40 sentences per dataset found a 7.5% annotation error rate, where models identified entities missed by human annotators, indicating their potential to improve pre-annotation processes.
@@ 3.5 Package Implementation
To support real-world application, we created an open-source clinical IE tool, ”Kiwi,” featuring our LLaMA-based and BiomedBERT-based models. Kiwi, available at https://kiwi.clinicalnlp.org/, is Python-based and packaged as a Docker image for cross-platform compatibility. It includes a CPU version with BiomedBERT and a GPU version with both BiomedBERT and LLaMA models. Kiwi docker image supports batch processing, and an online demo with a GUI is also provided on the website for testing purposes (see Figure 2 for interface screenshots).
@@ 4 Discussion
However, it’s important to note that while the performance gains are significant, they are not uniform across all tasks and datasets. When abundant annotated data are available at a healthcare system (e.g., UTP), the performance improvements of LLaMA are modest when compared with the BiomedBERT model (0.932 vs. 0.921 in F1 for NER and 0.937 vs 0.898 in F1 for RE), suggesting that deep learning models like BERT remain competitive in this scenario. Considering the low cost and high throughput of BERT models, we would recommend BERT models over LLMs when substantial annotated training data is already available for the IE task. However, when annotated training data is limited to a healthcare system (e.g., the case of i2b2), LLMs could be considered first, as they show better generalizability, indicating less annotation is needed to build the IE system. This highlights the need for careful consideration of the specific use case and available data when selecting a model for deployment.
@@
Computing Resources: The comparative analysis of computational requirements reveals important considerations for both inference and training. For inference deployment scenarios: (1) In high-resource settings with multiple NVIDIA A100 GPUs or better, LLaMA-3-70B offers the best performance; (2) In balanced settings with a single A100 GPU or equivalent, LLaMA-3-8B provides an attractive middle ground, achieving significant performance improvements; (3) For resource-constrained settings, BiomedBERT remains viable, requiring less resources while maintaining competitive performance on in-domain datasets. For training considerations, institutions have several options: (1) Using parameter-efficient tuning on larger LLaMA models (70B) demands significant resources but achieves the best performance; (2) Training smaller LLaMA models (7B/8B/13B) offers a balanced approach with competitive results; (3) Traditional fine-tuning of BiomedBERT provides the most resource-efficient though with lower performance on cross-institutional data. For institutions with extensive computational resources, pre-training domain-specific models like Me-LLaMA represents another option, though this requires more, making it practical only for large-scale research environments 45. These trade-offs underscore the importance of carefully matching computational capabilities with specific use case requirements.
Another significant contribution of this study is to release the Kiwi package. To the best of our knowledge, Kiwi is the first open-source clinical IE system based on LLMs. Kiwi includes both our instruction-tuned LLaMA-based and BiomedBERT-based models, available in both CPU and GPU versions to accommodate the diverse computational resources of different healthcare institutions. By offering different versions of models, Kiwi enables institutions to select the most appropriate model based on their specific needs, available resources, and use cases according to our suggestions. As a pioneering LLM-based clinical IE system, Kiwi represents a significant milestone in the application of generative AI to clinical IE. By making our system publicly available, we aim to accelerate research in LLM-based clinical IE by providing researchers with a ready-to-use framework. This will enable the medical NLP community to use our models directly, customize models upon our work, and develop more advanced clinical IE solutions.
@@ Acknowledgments
BiomedBERT 0.791 0.805 0.798 0.887 0.905 0.896 0.649 0.766 0.703 0.730 0.849 0.785
Exact and relax performance by LLaMA-2, LLaMA-3, and BiomedBERT on NER and RE across four datasets.
@@
BiomedBERT 0.761 0.806 0.783 0.847 0.918 0.881 0.589 0.686 0.634 0.692 0.793 0.739
Exact and relax performance by LLaMA-3-70B, and BiomedBERT on NER and RE across four datasets solely using MIMIC-III dataset for finetuning.

### v4

@@ Methods:
We investigated two fundamental clinical IE tasks: Named Entity Recognition (NER) and Relation Extraction (RE). Using 1,588 clinical notes from four data sources including UT Physicians (UTP), MTSamples, MIMIC-III, and i2b2, we developed a comprehensive annotated corpus covering 4 main clinical entities (problems, tests, medications, and other treatments) and 16 modifiers (e.g., negation, certainty). We instruction-tuned LLaMA-2 and LLaMA-3 for both tasks, comparing their performance, generalizability, computational resources, and throughput to BERT.
@@ Results:
LLaMA models consistently outperformed BERT across datasets and tasks. With sufficient training data (e.g., UTP), LLaMA showed marginal improvements (1% on NER, 1.5-3.7% on RE) over BERT; for limited training data (e.g., MTSamples, MIMIC-III), the improvements were greater. On the unseen i2b2 dataset, LLaMA-3-70B outperformed BERT by over 7% (F1) on NER and 4% (F1) on RE. However, LLaMA models required significantly more computing resources (i.e., memory and GPUs) and ran up to 28 times slower than BERT. To facilitate LLM use in clinical IE, we implemented ”Kiwi,” a clinical IE package featuring both BERT and LLaMA models and made it available at https://kiwi.clinicalnlp.org/.
@@ Conclusion:
This study is among the first to develop and evaluate a comprehensive clinical IE system using open-source LLMs. Results indicate that LLaMA models outperform BERT for clinical NER and RE but with higher computational costs and lower throughputs. These findings highlight that choosing between LLMs and traditional deep learning methods for clinical IE applications should remain task-specific, taking into account both performance metrics and practical considerations such as available computing resources and the intended use case scenarios.
@@ 1 Introduction
To address these challenges, we conducted an in-depth study to systematically compare LLM-based and BERT-based models on comprehensive clinical NER and RE tasks using a variety of evaluation metrics. Our study makes several key contributions to the clinical IE field. First, we developed one of the largest and most comprehensive cross-institutional annotated corpora for clinical NER and RE, comprising 1,588 clinical notes from multiple institutions. This corpus encompasses four primary clinical entities and 16 corresponding modifiers, providing a robust foundation for advancing clinical information extraction. Using the developed corpus, we systematically evaluated the performance of instruction-tuned LLaMA-3 models, demonstrating their superiority over previous state-of-the-art BERT-based models in clinical NER and RE tasks, particularly in low-resource and cross-institutional settings, with F1 score gains of up to 7% for NER and 4% for RE. Beyond performance metrics, we compared LLM-based and BERT-based clinical IE models across additional dimensions, including speed/throughput, computational resource requirements, and energy efficiency, which reveals some constraints of LLM-based IE models on real-world applications. Based on these findings, we provided detailed recommendations for model selection across diverse clinical NER and RE scenarios, bridging the gap between AI research and practical clinical use. Furthermore, we introduce Kiwi, one of the first publicly available LLM-based clinical IE systems. Kiwi provides both BERT-based and LLaMA-based models, allowing users to select the most appropriate tools based on their specific needs, computational resources, and task requirements, and it is freely accessible at https://kiwi.clinicalnlp.org/.
@@ 2.1 Task
Our study includes two clinical IE tasks: Named Entity Recognition (NER) and Relation Extraction (RE). NER identifies entity types and boundaries in text 33, while RE identifies relationships between entity pairs 34. We instruction-tuned LLaMA-2 and LLaMA-3 models and compared them to fine-tuned BERT. We focused on extracting four main entities—Medical Problems, Tests, Drugs and other Treatments—along with 16 modifiers (e.g., Severity, Temporal, Dosage). In total, there are 21 possible relations between main entities and modifiers, as detailed in Table 1. Annotation methods and guidelines are available in the Supplementary File and GitHub repository.
@@ 2.3.2 Bidirectional Encoder Representations from Transformers (BERT)
We fine-tuned BioClinicalBERT for NER and BiomedBERT for RE tasks 41; 18 for NER and RE tasks. Based on our preliminary experiments, these two distinct BERT models demonstrated superior performance for their respective tasks, motivating our decision to use them separately. For the NER model, we employed a learning rate of 5e-5 over 20 epochs with a batch size of 4. The RE model was trained using a learning rate of 1e-6 over 20 epochs with a batch size of 64.
@@ 2.5.1 Performance and significance test
We evaluate RE and NER performance using Precision (P), Recall (R), and F1 scores under exact and relaxed match criteria. Exact matches require entity types, boundaries, and relations to align perfectly, while relaxed matches require only entity type and relation matches with overlapping boundaries. Significance tests assess improvements over BERT, calculating p-values via a one-sided Wilcoxon rank-sum test with bootstrapping (sample size equal to test size, 1000 repetitions, 95% confidence interval).
@@ 2.5.4 Throughput
We assess each model’s throughput during fine-tuning and inference. For fine-tuning, we measure the total time to train and calculate seconds per note by dividing the total time by the number of notes in the dataset. For inference, we determine the time per note using the test dataset with a batch size of 100 for both LLaMA and BERT models.
@@ 3.1 Fine-tuning performance
Table 2 highlights the exact match performance of LLaMA-2, LLaMA-3, and BERT models, with full metrics in Supplementary Table S2. LLaMA models consistently outperformed BERT, particularly with limited fine-tuning data. With sufficient data (e.g., UTP dataset), LLaMA models show minor improvements of  1% on NER and 1.5-3.7% on RE. However, with less data (e.g., MTSamples, MIMIC-III), LLaMA models show greater performance improvement compared to BERT, with 3.7-4.5% higher F1 scores on NER and up to 12.5% higher on RE.
@@
BERT 0.898 0.670 0.808 0.703
Table 2: Exact F1 scores of LLaMA-2, LLaMA-3, and BERT on NER and RE across three datasets and cross-institution generalizability test results on unseen dataset (i2b2). All results of LLaMA models are statistically significant compared to BERT based on the one-sided Wilcoxon rank-sum test.
To explore the performance of LLaMA and BERT in low-resource settings, we conducted dataset-specific fine-tuning by training two models—LLaMA-3-70B and BERT—separately on the MTSamples and MIMIC-III datasets. Each model was fine-tuned exclusively on either MTSamples or MIMIC-III, then evaluated across all four test sets. Based on the results in Table 2, where LLaMA-3-70B demonstrated superior performance across datasets, we prioritized using this model for these experiments. Our findings show that LLaMA-3-70B consistently outperformed BERT across datasets in these low-resource scenarios, as shown in Supplementary Tables S3 and S4.
@@ 3.3 Computational resources and throughput
We compared the computational resources and throughput of BERT, LLaMA-2, and LLaMA-3 using NVIDIA A100 80GB GPUs (Tables 3 and 4). Fine-tuning was conducted with 1 GPU for BERT and the LLaMA. For inference, BERT and the 7B, 8B, and 13B LLaMA models used 1 GPU, whereas the 70B models required 2 GPUs. Results show notable trade-offs between performance and computational efficiency across model sizes.
@@ 3.3.1 Memory usage
Memory requirements increase substantially with model size. The 70B versions of LLaMA-2 and LLaMA-3 required 58-66 GB for training and 146 GB for inference, while BERT used only 21.3 GB and 65.7 GB, respectively. The 7B and 8B LLaMA models showed comparable memory efficiency to BERT, with training memory usage between 12.7-32.4 GB. Notably, during training, LLaMA models utilize Low-Rank Adaptation (LoRA), which reduces memory requirements compared to inference, where LoRA is not applied. This difference makes the memory needed for fine-tuning significantly lower than that required for inference. Additionally, these memory usages are heavily influenced by batch sizes; the current setup is optimized for best model performance.
@@ 3.3.2 Energy consumption and carbon emission
Energy consumption varied widely across models, with larger models requiring substantially more resources. LLaMA-70B models were the most energy-intensive during training, using 101-125 kWh and producing 39-49 kg of CO2). Smaller LLaMA models and BERT were more efficient, using 13-28 kWh and 5-11 kg of CO2). For inference, BERT was highly efficient at 0.03 kWh, while LLaMA models ranged from 0.14 kWh (LLaMA-3-8B) to 0.86 kWh (LLaMA-2-70B).
@@ 3.3.3 Training and Inference Throughput
Training and inference throughputs varied substantially across models and datasets as shown in table 4. LLaMA-3-8B had the fastest training time at 38.6 GPU hours, while BERT required 96.8 hours. BERT’s longer training time can be attributed to its need for 20 training epochs, compared to only 2 epochs for LLaMA models. This difference is reflected in GPU hours per epoch: BERT stands out with a particularly low requirement of 4.84 GPU hours per epoch, whereas LLaMA2-7B and LLaMA3-8B require 20.85 and 19.3, respectively. However, LLaMA-70B models took much longer, requiring 273.5-320.2 hours. Dataset characteristics such as note length impacted the training time; for example, MIMIC-III (2,106 tokens) took significantly more time per note than MTSamples (603 tokens). BERT outperformed all LLaMA models in inference, processing notes 5-28 times faster. Among LLaMA models, LLaMA-3-8B was quickest, while the 70B models were slowest. These findings highlight a trade-off: larger LLaMA models offer higher accuracy but are slower, while BERT is more efficient for time-sensitive applications.
@@ 3.4 Error analysis
Automatic and manual error analyses on LLaMA and BERT outputs across four datasets revealed that LLaMA-70B models have fewer False Negatives (FN) but slightly higher False Positives (FP) than smaller models, indicating a precision-recall trade-off with model size. BERT, while showing higher FN and FP rates, has fewer boundary errors, suggesting better accuracy in detecting text boundaries. Type errors were consistently low for all models, confirming effective classification once detected. Manual analysis of 40 sentences per dataset found a 7.5% annotation error rate, where models identified entities missed by human annotators, indicating their potential to improve pre-annotation processes.
@@ 3.5 Package Implementation
To support real-world application, we created an open-source clinical IE tool, ”Kiwi,” featuring our LLaMA-based and BERT-based models. Kiwi, available at https://kiwi.clinicalnlp.org/, is Python-based and packaged as a Docker image for cross-platform compatibility. It includes a CPU version with BERT and a GPU version with both BERT and LLaMA models. Kiwi docker image supports batch processing, and an online demo with a GUI is also provided on the website for testing purposes (see Figure 2 for interface screenshots).
@@ 4 Discussion
However, it’s important to note that while the performance gains are significant, they are not uniform across all tasks and datasets. When abundant annotated data are available at a healthcare system (e.g., UTP), the performance improvements of LLaMA are modest when compared with the BERT model (0.932 vs. 0.921 in F1 for NER and 0.937 vs 0.898 in F1 for RE), suggesting that deep learning models like BERT remain competitive in this scenario. Considering the low cost and high throughput of BERT models, we would recommend BERT models over LLMs when substantial annotated training data is already available for the IE task. However, when annotated training data is limited to a healthcare system (e.g., the case of i2b2), LLMs could be considered first, as they show better generalizability, indicating less annotation is needed to build the IE system. This highlights the need for careful consideration of the specific use case and available data when selecting a model for deployment.
@@
Computing Resources: The comparative analysis of computational requirements reveals important considerations for both inference and training. For inference deployment scenarios: (1) In high-resource settings with multiple NVIDIA A100 GPUs or better, LLaMA-3-70B offers the best performance; (2) In balanced settings with a single A100 GPU or equivalent, LLaMA-3-8B provides an attractive middle ground, achieving significant performance improvements; (3) For resource-constrained settings, BERT remains viable, requiring less resources while maintaining competitive performance on in-domain datasets. For training considerations, institutions have several options: (1) Using parameter-efficient tuning on larger LLaMA models (70B) demands significant resources but achieves the best performance; (2) Training smaller LLaMA models (7B/8B/13B) offers a balanced approach with competitive results; (3) Traditional fine-tuning of BERT provides the most resource-efficient though with lower performance on cross-institutional data. For institutions with extensive computational resources, pre-training domain-specific models like Me-LLaMA represents another option, though this requires more, making it practical only for large-scale research environments 45. These trade-offs underscore the importance of carefully matching computational capabilities with specific use case requirements.
Based on these findings, we made some recommendations for model choice for different clinical IE tasks in different resource settings, summarized in Figure 3.
Figure 3: Recommendations for model choice for different tasks in different resource settings.
Another significant contribution of this study is to release the Kiwi package. To the best of our knowledge, Kiwi is the first open-source clinical IE system based on LLMs. Kiwi includes both our instruction-tuned LLaMA-based and BERT-based models, available in both CPU and GPU versions to accommodate the diverse computational resources of different healthcare institutions. By offering different versions of models, Kiwi enables institutions to select the most appropriate model based on their specific needs, available resources, and use cases according to our suggestions. As a pioneering LLM-based clinical IE system, Kiwi represents a significant milestone in the application of generative AI to clinical IE. By making our system publicly available, we aim to accelerate research in LLM-based clinical IE by providing researchers with a ready-to-use framework. This will enable the medical NLP community to use our models directly, customize models upon our work, and develop more advanced clinical IE solutions.
@@ Acknowledgments
BERT 0.791 0.805 0.798 0.887 0.905 0.896 0.649 0.766 0.703 0.730 0.849 0.785
Exact and relax performance by LLaMA-2, LLaMA-3, and BERT on NER and RE across four datasets.
@@
BERT 0.761 0.806 0.783 0.847 0.918 0.881 0.589 0.686 0.634 0.692 0.793 0.739
Exact and relax performance by LLaMA-3-70B, and BER on NER and RE across four datasets solely using MIMIC-III dataset for finetuning.